40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 8 04:13:12 2012, kiwamu, Update, LSC, update on the locking activity 
    Reply  Fri Mar 9 13:34:13 2012, keiko, Update, LSC, update on the locking activity 
       Reply  Sat Mar 10 02:00:03 2012, keiko, Update, LSC, update on the locking activity 
Message ID: 6398     Entry time: Sat Mar 10 02:00:03 2012     In reply to: 6393
Author: keiko 
Type: Update 
Category: LSC 
Subject: update on the locking activity 

ITMX and ITMY balance for the MICH excitation (lockin) is adjusted again. Now it's ITMx = -0.992, ITMy = 1 for MICH (lockin output matrix values).

RA: what were the old values? Does this change make any difference for the signal mixing noticed before?

ELOG V3.1.3-